OpenStudy (anonymous):

find the integral of (sin2x)/(cosx) dx without using substitution.

zepdrix (zepdrix):

is the 2 a 2, or an exponent? :U

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{\sin2x}{\cos x}dx\]Like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

double angle?

OpenStudy (anonymous):

i got as far as this. \[\int\limits \frac{ 2sinxcosx }{ cosx } dx\]

OpenStudy (kirbykirby):

can you see that \(\cos(x)\) cancels out :)